# Error correction of transversal controlled-NOT gates for scalable surface code computation

### v1

@@ III A scalable transversal CNOT
To this end, first consider $r,\;m\ll W$ for which, following the arguments from the earlier paragraph, we find that a decoder can fail to correct a $Z$-error string $E$ of length $O(mr)$ on $C_{1}$ at the beginning of the window i.e., the first tCNOT. So clearly, $r,m\ll W$ is not sufficient to prevent a long-lived $Z$ error. On the other hand, setting $mr=W$ is sufficient to prevent $Z$ errors at the beginning of the window from surviving for a long time. As expected, this condition cannot be satisfied if both $m$ and $r$ are constants. However, it can be satisfied by setting $r=O(1)$ but $m=O(W)$, although, in this case the decoding volume increases exponentially with $W$ or equivalently exponentially with $d$, leading to an undesirable slowdown in decoding. Clearly, the simplest way to achieve the condition $mr=W$ without increasing the decoding complexity exponentially with $poly(d)$ is by choosing $r=O(W)$ and $m=O(1)$.
Hence we find that even in the expanded window approach $r=O(W)$ is desirable. Moreover, a reason why $r=W$ and $m=1$, which effectively means that each tCNOT is decoded independently, is desirable is because it allows true benchmarking of a tCNOT gate. When $m\neq 1$ the internal structure of the circuit determines how errors spread between different code blocks in the window size $m$, thus making decoding circuit-dependent. Consequently, for efficient decoders, no absolute algorithm-independent performance benchmark may be guaranteed if $m\neq 1$, further enforcing the utility of individual tCNOT decoding enabled by $r=W$.
@@
Having addressed the effectiveness of $r=W$, we now discuss three decoding procedures that can be used to decode an individual tCNOT. Specifically we use the simplifying assumption of $W=d$ since the measurement errors in our numerically simulated noise model are roughly as likely as data qubit errors.
@@ IV Decoders for a transversal CNOT
Figure 4: Thresholds and logical failure rates under circuit-level depolarising Pauli noise for SCQMs and logical CNOTs. (left) HUF decoder performance for a 2SCQM and tCNOT. (centre) MWPM decoder performance for a 2SCQM and lattice-surgery (LS) CNOT. (right) single-update and ordered MWPM decoder performance for a tCNOT. Operations are plotted separately to prevent overcrowding. Thresholds are found by finite size scaling close to $p_{t}$ (translucent lines).
@@
For our decoding analysis, we introduce the notion of dynamic and static frames for the check stabilizers. In what we term the dynamic frame, each $X$ check used for correcting $Z$ errors evolves through the tCNOT as
@@
This reflects the internal evolution of the checks caused by the tCNOT. $\mathbf{S}^{i}_{CX}=S^{i}_{CX},\forall i\leq d$ and $\mathbf{S}_{TX}=S_{TX}$, i.e. other checks remain unchanged in the dynamic frame. In contrast, we define a static frame, where all checks are left in their original states.
For any given frame, we can define the $X$-decoding graph of the system $G$ as laid out in Sec. II. For convenience in the chosen frame, we break up $G_{X}$ into subgraphs $G_{CX}$ and $G_{TX}$, where $G_{CX}$ ($G_{TX}$) contains the frame-defined nodes related to the checks of $C$ $(T)$. We term $G_{CX}$ a dependent subgraph as its checks change due to the tCNOT in the dynamic frame. From a complementary perspective, the control can be regarded as dependent because $Z$ errors are copied over to $C$ by the tCNOT. Unlike $G_{CX}$, the nodes of $G_{TX}$ are unaffected by the tCNOT; it is thus termed an independent subgraph. Note that in experiment, we always only measure checks in the static frame. Checks of the dynamic frame can be inferred by combining the outcomes of the static $X$-checks at the cost of making the inferred dynamic checks more unreliable than the measured static checks in the presence of measurement errors.
@@ IV.1 Single-update decoder
In the single-update decoding strategy, we operate in the dynamic frame. Post tCNOT, since we update the $S_{CX}$ check measurement outcomes according to Eqn. 1, the detectors $v_{CX}^{{i}}=\mathbf{S}_{CX}^{i-1}\oplus\mathbf{S}_{CX}^{i}\quad\forall i\in\{1,2,...,2d\}$ correctly track checks through the tCNOT. In terms of the measured static checks, we now have
$v_{CX}^{{d}}=S_{CX}^{d-1}\oplus S_{CX}^{d},$ (2a)
$v_{CX}^{{d+1}}=S_{CX}^{d}\oplus S_{CX}^{d+1}\oplus S_{TX}^{d+1},\text{ and}$ (2b)
$v_{CX}^{{d+2}}=S_{CX}^{d+1}\oplus S_{TX}^{d+1}\oplus S_{TX}^{d+2}\oplus S_{TX}^{d+2}$ (2c)
We can now apply MWPM to decode the defects created in the new dynamic-frame detector set, similar to that of an SCQM. This is explicitly detailed in Alg. 1. We illustrate key differences between the single-update decoder and an SCQM using an instance of a data qubit error $Z_{T,q}$ on a qubit $q$ in $T$ that creates defects on $T$ at round $k$:
  1. 1.
    If $k\leq d$: After round $d$, $Z_{T,q}$ is copied over to the control, creating $Z_{C,q}$. Since $\left[Z_{T,q}Z_{C,q},S_{CX}S_{TX}\right]=0$, no defect is created in the dynamic frame after the tCNOT despite the actual error being present. This behaviour differs from what happens when decoding in static frame, as will be evident from analysis in the following section.
  2. 2.
    If $k>d$: $Z_{T,q}$ is not copied over to $C$, but creates new defects on the updated $C$ due to the updated checks of the dynamic frame: $\left[Z_{T,q},\mathbf{S}_{CX}=S_{CX}S_{TX}\right]\neq 0$. These errors are not truly on the control, and the resulting ‘false defects’ simply mirror defect patterns on the target. Note that these false defects appear in addition to those created by errors on the target itself, resulting in an effective doubling of the defect rate post-tCNOT on $T$ that are required to be dealt with by a decoder.
We observe that use of the dynamic frame prevents propagated errors from creating defects on $C$, while creating false defects from errors that did not propagate. After independently decoding and correcting both subgraphs in the dynamic frame, we use knowledge of all the corrections applied to $T$ to apply a final logical update to $C$ and restore the code states.
To benchmark the performance of this single-update decoder, we compare the numerically calculated logical error rate for the tCNOT decoded in this fashion to that for two independent, disjoint SCQMs with for $2d$ rounds of check measurements (hereon referred to as a 2SCQM experiment). The results are shown in Fig. 4. In this and all circuit-level simulations with Pauli noise, two-qubit gate errors are uniformly chosen at random from $\{I,X,Y,Z\}^{\otimes 2}/\{I\otimes I\}$ at a rate $p$ while assuming no single-qubit gate, initialization or measurement errors  [fowler_surface_2012]. As expected, due to the larger defect population on $C$, this procedure results in a reduction in threshold (${p_{t}}(\text{tCNOT})=0.6\%$) compared to a 2SCQM. We note that the single-update decoding strategy was proposed in Ref. [dennis_topological_2002] and previously used in Ref. [beverland_cost_2021].
@@ IV.2 Combined hypergraph decoder
The dynamic frame, while natively following the tCNOT induced check evolution, increases the number of errors on the dependent subgraph. It is thus natural to attempt to use the static frame where possible. As explained in Appendix A, it is possible to use a combination of the static and dynamic frames to define a set of detectors with minimal overlap between $C$ and $T$ while attempting to respect the tCNOT evolution. We call this the hybrid frame. In this frame, the tCNOT creates non-decomposable hyperedges in the decoding graph, i.e. a single independent error mechanism causes three or more defects. This has been shown in Refs.  [cain_correlated_2024, zhou_algorithmic_2024]. A naive matching decoder cannot successfully decode hyperedges; we thus use a hypergraph decoder, specifically the hyper-graph union find (HUF) decoder [delfosse_toward_2021, delfosse_almost-linear_2021] implemented in the MWPF package [wu_wuyue16pkugmailcom_mwpf_nodate].
[…]
Hypergraph decoding presents a potential path to decode and correct logical operations that induce hyperedges. However, as demonstrated above, time-efficient hypergraph strategies underperform with increased hypergraph complexity and have higher runtime overheads than their graph-based counterparts. Ideally, we would like an efficient decoding algorithm for the tCNOT that scales equal to or better than the equivalent decoder applied to a 2SCQM, while at the same time preserving the SCQM threshold. Our next decoding strategy, ordered decoding, achieves this goal.
@@ IV.3 Ordered decoder
Let us now describe the functioning of an ordered decoding strategy. This operates entirely in the static frame. As hinted at in the single update decoder, $Z$ errors occurring on the target pre-tCNOT - and only these $Z$ errors - are propagated to the control. If we can, to the best of our ability, fully identify such errors via decoding the target first, we then know exactly what defects they will create on the control at round $d+1$. Ordered decoding relies on exactly this principle: we first decode $T$ using the decoding graph $G_{TX}$ in the static frame. From the resulting solution, the decoder identifies error clusters on $T$ that occur before the tCNOT. Detectors in $G_{CX}$ in the static frame that correspond to the propagation of these identified errors from target to control at round $d+1$ are then flipped. This process changes the collection of apparent defects on the control. We finally simply decode and correct the control with the updated defect collection. A corresponding mechanism can also be applied to the $Z$-decoding graph with the control decoded before the target (see Alg. 2 for a full description).
An ordered decoding strategy results in the preservation of the threshold (see Fig. 4 with ${p_{t}}(\text{tCNOT})=1.03\%$ where a MWPM decoder is used for the individual decoding steps), with only a marginal bounded increase in logical error rates compared to an equivalent decoder on a 2SCQM (see App. C for further analysis). Note that ordered decoding doubles the decoding time as decoding $C$ for $Z$ errors can only begin after the target is decoded (and vice-versa for $X$ errors). This constant factor increase does not amplify the backlog problem.
@@ V Logical state teleportation
One of the primary uses of two-qubit gates in quantum algorithms will be for logical state teleportation, particularly for non-Clifford gates, such as in Fig. V. Here, we investigate an instance of such fault-tolerant logical state teleportation circuits, focusing on the teleportation step that may be implemented transversally or by joint-parity measurements. In a transversal teleportation circuit using a tCNOT, where a logical measurement is to be immediately implemented following the tCNOT, we find that, unlike the general unitary case, subsequent stabilizer measurement rounds between the tCNOT and the logical measurement on the target are not necessary to maintain code performance. We additionally compare a transversal teleportation strategy to that using joint measurements, an approach generally considered more suitable for planar architectures.
@@
Figure 5: Methods to perform fault-tolerant magic state state teleportation. (a) A logical state is teleported onto the original qubit using a CNOT, followed by a Clifford update. (b) In a lattice-surgery setting, the equivalent protocol in (a) can be optimized to reduce the number of joint parity measurements [litinski_game_2019].
@@ V.1 Hypergraph reduction for transversal measurements
We first study the transversal teleportation scheme in Fig. V(a), where a logical measurement of one of the surface codes immediately succeeds a tCNOT. We consider $d$ rounds of stabilizer measurements on both the control and the target before the tCNOT. Post tCNOT, since the logical $Z$ measurement on $T$ determines whether an $S$ gate is applied to the control, it is desirable that the logical $Z$ measurement is correctly read out.
We find that the logical measurement terminates the decoding graph and consequently reduces the hyperedges across the tCNOT gates in the hybrid frame (App. A) to edges. Therefore, MWPM decoders are naturally applicable in this scenario. To explain this phenomenon, we first focus on correction of the logical $Z$ measurement on the target. The logical $Z$ measurement is performed by measuring all the data qubits of $T$ in the $Z$ basis after the tCNOT. From these measurement results, we can extract both the $Z$ logical measurement outcome and an extra $(d+1)$-th round of $Z$ stabilizer measurements on the target from the product of measurement results on their corresponding support qubits.
Here, before the tCNOT, we have detectors $v^{i}_{CZ}=S^{i-1}_{CZ}\oplus S^{i}_{CZ}$ and $v^{i}_{TZ}=S^{i-1}_{TZ}\oplus S^{i}_{TZ}\quad\forall i\in\{1,2,...,d\}$. Across the tCNOT gate, we have an additional detector of the hybrid frame $v^{d+1}_{TZ}=S^{d}_{CZ}\oplus S^{d}_{TZ}\oplus S^{d+1}_{TZ}$. Crucially, unlike the tCNOT circuit we study in Sec IV, we do not include $d$ rounds of stabilizer measurement results on the control following the tCNOT during this correction. Therefore, there are no additional detectors $v^{i}_{CZ}$ used for $i>d$ after the tCNOT. Since each check is now involved in at most two detectors, this detector subset reduces hyperedges to edges. For example, a measurement error on $S^{d}_{CZ}$ flips only two detectors $v^{d}_{CZ}$ and $v^{d+1}_{CZ}$ instead of three as described in Appendix A. For correction of the logical $Z$ measurement on $T$ with an MWPM decoder, we obtain a threshold of 1.12% , as shown in Fig. 6.
We now discuss $X$ error correction on the control. During correction of the logical $Z$ measurement on $T$, we can also obtain a correction for pre-tCNOT control $X$ errors. We first apply this correction to $C$ before continuing to measure $d$ rounds of stabilizers, which can subsequently be treated as an SCQM. Fig. 6 shows the logical error rate in this case, which overlaps with that of the target logical $Z$ measurement.
We now move onto $Z$ error correction on $C$ using $X$ checks. Across the tCNOT gate, the hybrid frame detectors are $v^{d+1}_{CX}=S^{d}_{CX}\oplus S^{d}_{TX}\oplus S^{d+1}_{CX}$. Unlike the tCNOT circuit we study in Sec IV, we do not have $d$ post-tCNOT rounds of stabilizer measurement results on the target. Therefore, there are no additional detectors $v^{i}_{TX}$ for $i>d$ after the tCNOT. As a result, a measurement error on $S^{d}_{TX}$ only flips two detectors $v^{d}_{TX}$ and $v^{d+1}_{CX}$ instead of three. Once again, there are no hyperedges, and we can decode this system using MWPM. We thus show that graph-based decoding and its inherent logical error rates and thresholds are maintained when specific detectors are not present, as is in the case of transversal logical measurements immediately after a tCNOT.

### v2

@@ III A scalable transversal CNOT
To this end, first consider $r,\;m\ll W$ for which, following previous arguments, we find that a decoder can fail to correct a $Z$-error string $\mathcal{E}$ of length $O(mr)$ on $C_{1}$ at the beginning of the window i.e., the first tCNOT. Thus, $r,m\ll W$ is not sufficient to prevent a long-lived $Z$ error. On the other hand, setting $mr=W$ is sufficient to prevent $Z$ errors at the beginning of the window from surviving for a long time. As expected, this condition cannot be satisfied if both $m,r=O(1)$. It can instead be satisfied by setting $r=O(1)$ and $m=O(W)$. In this case the decoding volume increases exponentially with $m$, leading to an undesirable slowdown in decoding. Clearly, the simplest way to achieve $mr=W$ without increasing the decoding volume exponentially is by choosing $r=O(W)$ and $m=O(1)$. These conditions are summarized in Table I.
Hence we find that even in the expanded window approach $r=O(W)$ is desirable. Moreover, when $r=W$ and $m=1$, each tCNOT gadget is effectively decoded independently. The advantage of these modular gadgets is that one can benchmark their individual failure rates and then estimate algorithmic performance for a wide variety of algorithms. For increasingly larger gadgets, the internal structure of the circuit determines how errors spread between different code blocks, complicating later circuit analysis. To predict algorithmic performance, a defined set of gadgets would need to be benchmarked and the errors carefully composed.
@@
Having addressed the effectiveness of $r=W$, we now discuss three decoding procedures that can be used to decode an individual tCNOT. Specifically we use the simplifying assumption of $W=d$ since measurement errors in our numerically simulated noise model are roughly as likely as data qubit errors.
@@ IV Decoders for a transversal CNOT
Figure 4: Thresholds and logical failure rates under circuit-level depolarising Pauli noise for SCQMs and logical CNOTs. (left) HUF decoder performance for a 2SCQM and tCNOT. (centre) single-update and ordered MWPM decoder performance for a tCNOT. (right) MWPM decoder performance for a 2SCQM and lattice-surgery (LS) CNOT. Operations are plotted separately to prevent overcrowding. Thresholds are found by finite size scaling close to $p_{t}$ (translucent lines).
@@
For our decoding analysis, we introduce the unifying perspective of check frames for SCs in the tCNOT. In what we term the dynamic frame, each $X$ check used for correcting $Z$ errors evolves through the tCNOT as
@@
This reflects the internal evolution of the checks caused by the tCNOT. $\mathbf{S}^{i}_{CX}=S^{i}_{CX},\forall i\leq d$ and $\mathbf{S}_{TX}=S_{TX}$, i.e. other $X$-checks remain unchanged in the dynamic frame. In contrast, we define a static frame, where all checks are left in their original values.
For any given frame, we can define the $X$-decoding graph of the system $G$ as laid out in Sec. II. For convenience in the chosen frame, we break up $G_{X}$ into subgraphs $G_{CX}$ and $G_{TX}$, where $G_{CX}$ ($G_{TX}$) contains the frame-defined nodes related to the checks of $C$ $(T)$. We term $G_{CX}$ a dependent subgraph since its checks change due to the tCNOT in the dynamic frame. From a complementary perspective, the control can be regarded as dependent because $Z$ errors are copied over to $C$ by the tCNOT. Unlike $G_{CX}$, the nodes of $G_{TX}$ are unaffected by the tCNOT; it is thus termed an independent subgraph.
Note that in experiment, we always only measure checks in the static frame. Checks of the dynamic frame can be inferred by combining the outcomes of the static $X$-checks at the cost of making the inferred dynamic checks more unreliable than the measured static checks in the presence of measurement errors.
@@ IV.1 Single-update decoder
In the single-update decoding strategy, we operate in the dynamic frame. Post-tCNOT, since we update the $S_{CX}$ check measurement outcomes according to Eqn. 1, the detectors $v_{CX}^{{i}}=\mathbf{S}_{CX}^{i-1}\oplus\mathbf{S}_{CX}^{i}$ correctly track checks through the tCNOT. We can now apply MWPM to decode the defects created in the new dynamic-frame detector set, similar to that of an SCQM. This is explicitly detailed in Alg. 1. Note that the use of the dynamic frame doubles the defect rate on $G_{CX}$ post-tCNOT. We provide examples of handling of relevant errors by the single-update decoder in Appendix A.
To benchmark the performance of the single-update decoder, we compare the numerically calculated logical error rate for the tCNOT decoded in this fashion using MWPM, to that of two independent, disjoint SCQMs for $2d$ rounds of check measurements (hereon referred to as a 2SCQM experiment). The results are shown in Fig. 4 centre and right. In this and all circuit-level simulations with Pauli noise, two-qubit gate errors are uniformly chosen at random from $\{I,X,Y,Z\}^{\otimes 2}/\{I\otimes I\}$ at a rate $p$ while assuming no single-qubit gate, initialization or measurement errors [fowler_surface_2012] (see App I). As expected, due to the larger defect population on $G_{CX}$, this procedure results in a reduction in tCNOT threshold (${p_{t}}=0.6\%$) compared to a 2SCQM (${p_{t}}=1.04\%$). We note that the single-update decoding strategy was proposed in Ref. [dennis_topological_2002] and previously used in Ref. [beverland_cost_2021].
@@ IV.2 Combined hypergraph decoder
We now move onto our second decoding strategy. The dynamic frame, while natively following the tCNOT induced check evolution, increases the number of errors on the dependent subgraph. It is thus natural to attempt to use the static frame where possible. As explained in Appendix B, it is possible to only use the dynamic frame at the tCNOT round, with all other rounds in the static frame to define a set of detectors. We call this the hybrid frame. In this frame, the tCNOT creates non-decomposable hyperedges in the decoding graph, i.e. a single independent error mechanism creates three or more defects. This has been shown in Refs.  [cain_correlated_2024, zhou_algorithmic_2024, hetenyi_creating_2024]. A naive matching decoder cannot successfully decode hyperedges; we thus use a hypergraph decoder, specifically the hypergraph union find (HUF) decoder [delfosse_toward_2021, delfosse_almost-linear_2021, wu_wuyue16pkugmailcom_mwpf_nodate].
We compare the performance of the tCNOT decoded with HUF to that of 2SCQM in Fig. 4 (left). Interestingly, we see a slight reduction in threshold for the tCNOT (${p_{t}}=0.78\%$) compared to the 2SCQM experiment (${p_{t}}=0.89\%$). This may be attributed to the locally – as opposed to globally – optimal corrections found by the HUF algorithm in combination with the increased complexity of hyperedges in the decoding graph for the tCNOT versus the 2SCQM.
Hypergraph decoding presents a potential path to decode and correct logical operations that induce hyperedges. However, as seen above, time-efficient hypergraph strategies underperform with increased hypergraph complexity and have higher runtime overheads than their graph-based counterparts. Ideally, we would like an efficient decoding algorithm for the tCNOT that scales equal to or better than the equivalent decoder applied to a 2SCQM, while at the same time preserving the SCQM threshold. Our next decoding strategy, ordered decoding, achieves this goal.
@@ IV.3 Ordered decoder
Let us now describe an ordered decoding strategy. This operates entirely in the static frame. We know that $Z$ errors occurring on the target pre-tCNOT – and only these $Z$ errors – are propagated to the control. If we can, to the best of our ability, fully identify such errors via decoding the target first, we then know exactly what defects they will create on the control at round $d+1$.
Ordered decoding relies on exactly this principle: we first decode $T$ using $G_{TX}$ in the static frame. From the resulting solution, the decoder identifies error clusters on $T$ that occur before the tCNOT. Detectors in $G_{CX}$ in the static frame that correspond to the propagation of these identified errors from target to control at round $d+1$ are then flipped, keeping track of the updated logical status. This process changes the collection of apparent defects on the control. We finally simply decode and correct the control with the updated defect collection. A corresponding mechanism can also be applied to the $Z$-decoding graph with the control decoded before the target (see Alg. 2 for a full description).
An ordered decoding strategy on a tCNOT results in the preservation of the SCQM threshold (see Fig. 4 (centre) with ${p_{t}}=1.03\%$), with only a marginal bounded increase in logical error rates compared to an equivalent decoder on a 2SCQM (see App. D for further analysis). Note that ordered decoding doubles the decoding time as decoding $C$ for $Z$ errors can only begin after the target is decoded (and vice-versa for $X$ errors). This constant factor increase does not amplify the backlog problem.
@@ V Logical state teleportation
Figure 5: Methods to perform a $T$ gate using fault-tolerant magic state state teleportation. (a) A logical state is teleported onto the original qubit using a CNOT, followed by a Clifford update. (b) In a lattice-surgery setting, the equivalent protocol in (a) can be optimized to reduce the number of joint parity measurements [litinski_game_2019].
One of the primary uses of two-qubit gates in quantum algorithms will be for logical state teleportation, particularly for non-Clifford gates, such as in Fig. V. Here, we investigate these fault-tolerant teleportation circuits, focusing on the teleportation step that may be implemented transversally or by joint-parity measurements. In a transversal teleportation circuit using a tCNOT, where a logical measurement is to be immediately implemented following the tCNOT, we find that, unlike the general unitary case, subsequent stabilizer measurement rounds between the tCNOT and the logical measurement on the target are not necessary to maintain code performance. We additionally compare a transversal teleportation strategy to that using joint measurements, an approach generally considered more suitable for planar architectures.
@@ V.1 Hypergraph reduction for transversal measurements
We first study the transversal teleportation scheme in Fig. V(a), where a logical measurement of one of the surface codes immediately succeeds a tCNOT. We consider $d$ rounds of stabilizer measurements on both the control and the target before the tCNOT. Post-tCNOT, since the logical $Z$ measurement on $T$ determines whether an $S$ gate is applied to the control, it is desirable that the logical $Z$ measurement is correctly read out. We find that the logical measurement terminates the decoding graph and reduces the hybrid frame hyperedges across the tCNOT to edges. Therefore, MWPM decoders are naturally applicable in this scenario.
@@
We first focus on correction of the logical $Z$ measurement on the target. From the measurement results of all data qubits of $T$ in the $Z$ basis after the tCNOT, we can extract both the $Z$ logical measurement outcome and an extra $(d+1)$-th round of $Z$ stabilizer measurements from products of measurement results. We use detectors of the hybrid frame (App. B). Crucially, unlike the tCNOT circuit we study in Sec IV, we do not include $d$ rounds of stabilizer measurement results on the control following the tCNOT during this correction. Therefore, there are no additional detectors $v^{i}_{CZ}$ used for $i>d$ after the tCNOT. A measurement error on $S^{d}_{CZ}$ thus flips only two detectors $v^{d}_{CZ}$ and $v^{d+1}_{CZ}$ instead of three as described in Appendix B. Since each check is now involved in at most two detectors, this detector subset reduces hyperedges to edges. For correction of the logical $Z$ measurement on $T$ with an MWPM decoder, we obtain $p_{t}=1.12\%$, as shown in Fig. 6.
We now move on to $X$ error correction on $C$. During correction of the logical $Z$ measurement on $T$, we also obtain a correction for pre-tCNOT control $X$ errors. We first apply this correction to $C$ in an ordered decoding manner before continuing to measure $d$ rounds of stabilizers, which can subsequently be treated as an SCQM. Fig. 6 shows the logical error rate in this case, which overlaps with that of the target logical $Z$ measurement.
For $Z$ error correction on $C$ using $X$ checks, we similarly find no hyperedges. Thus, graph-based decoding and its inherent logical error rates and thresholds are maintained when specific detectors are not present, as is in the case of transversal logical measurements immediately after a tCNOT.